OpenStudy (anonymous):

extremely easy, need assistance? Medal! so I'm in college algebra now, and I seem to have forgotten some basics. Can someone help me solve for y in this equation x=9y+7/7y-2

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi!

OpenStudy (misty1212):

this is not really that easy so it is easy to forget

OpenStudy (anonymous):

lol okay, thanks for the reassurance.

OpenStudy (misty1212):

\[x=\frac{9y+7}{7y-2}\] right?

OpenStudy (anonymous):

yup!

OpenStudy (misty1212):

first multiply both sides by \(7y-2\) to clear the fraction

OpenStudy (anonymous):

should you always do the denominator first?

OpenStudy (misty1212):

you get \[x(7y-2)=9y+7\] then multiply out on the right

OpenStudy (misty1212):

you gotta get rid of the denominator to solve, yes

OpenStudy (anonymous):

okay, and that's what I got also.

OpenStudy (misty1212):

when you multiply out you get \[7xy-2x=9y+7\]

OpenStudy (anonymous):

so 7yx-2x=9y+7

OpenStudy (misty1212):

right

OpenStudy (misty1212):

now put all the stuff with \(y\) on one side of the equal sign, (say the left) everything else on the other

OpenStudy (anonymous):

okay, one second

OpenStudy (anonymous):

I got 5x-7=9y/y

OpenStudy (misty1212):

no not quite

OpenStudy (anonymous):

how do i get the y alone then?

OpenStudy (anonymous):

that's where I always mess it up.

OpenStudy (misty1212):

we are not there yet, hold on

OpenStudy (misty1212):

we are good to here \[7xy-2x=9y+7\] now you want all the y terms on the left, subtract \(9y\) from both sides, which does not really mean "subtract" it just means "write it" \]7xy-9y-2x=7\]

OpenStudy (misty1212):

oops \[7xy-9y-2x=7\] that's better

OpenStudy (anonymous):

oh... I went about it the wrong way...

OpenStudy (misty1212):

then add \(2x\) to both sides what do you get?

OpenStudy (anonymous):

7yx-9y=2x+7

OpenStudy (misty1212):

ok good so we have all the y terms on the left, all else on the right that is correct because we want to get y by itself now factor out the y from the left do you know what i mean by that?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

y(7x-9)=2x+7

OpenStudy (misty1212):

good now one step and you are done, namely divide by \(7x-9\) to get \(y\)

OpenStudy (anonymous):

and then divide by the 7x-9

OpenStudy (misty1212):

yup `

OpenStudy (anonymous):

okay, got it. Thank you so much. lol that was much more difficult than I remember learning in 6th grade. I caught on to bad math habits and it shows in how I try to do things now ):

OpenStudy (misty1212):

not terrifically hard, but not really that easy take several steps

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Thanks again! have a great day!

OpenStudy (misty1212):

you too